solve by substitution 3x+y= -3
y=x+5

Answer by stanbon(75887)

solve by substitution
3x+y = -3
y = x+5
---------
substitute for "y" and solve for "x":
3x + x+5 = -3
4x = -8
x = -2
-----
Since y = x+5, y = -2+5 = 3
==============================
Cheers,
Stan H.